OpenStudy (anonymous):

solve the exponential equation. Use a calculator to obtain a decimal approximation, correct to two decimal places, for the solution. 3e^x=29 I need help on this ASAP!!!!!!

OpenStudy (anonymous):

Its been a while but i belive you want to use ln so that you dont have x as an exponent..hope this helps

OpenStudy (anonymous):

@jigglypuff314 is there any way you could help me on this?????

OpenStudy (anonymous):

@texaschic101 may you please help me on this?????

jigglypuff314 (jigglypuff314):

I haven't done ln or log stuff in a long while lol idk... umm... divide both sides by 3 convert to ln... then plug into calculator.

OpenStudy (anonymous):

@jigglypuff314 so i would put ln29/3 into the calculator?

jigglypuff314 (jigglypuff314):

yep...

OpenStudy (anonymous):

@jigglypuff314 when i did that i didn't get the right answer :(

jigglypuff314 (jigglypuff314):

aww >,< sorry :/

OpenStudy (anonymous):

1) divide by 3 2) take the log

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=log%2829%2F3%29

OpenStudy (anonymous):

 @satellite73 this is what i have so far

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

\[3e^x=29\\ e^x=\frac{29}{3}\\ x=\ln(\frac{29}{3})\]

jigglypuff314 (jigglypuff314):

divide both sides by 3 you get e^x = 29/3 covert to log or ln so ln(29/3) = 2.27

OpenStudy (anonymous):

@satellite73 @jigglypuff314 thank you so much!!!!! i understand now.